OpenStudy (anonymous):

Can somebody walk me through how to solve this Algebra problem?? I'm stuck. "Solve the following systems of a circle and a line by finding the intersection point(s), if any, algebraically. 3. y=x-4 (x+2)^2+y^2=4

OpenStudy (campbell_st):

I just graphed the 2 equations... and they don't intersect...

OpenStudy (anonymous):

so now, it would be.... \[x ^{2}+4x+4+x ^{2}-8x+16=4\] before i combine like terms?

OpenStudy (campbell_st):

that's the correct process... but the resulting quadratic will have complex roots. so the lines don't intersect

OpenStudy (anonymous):

so should i just write that??? because i have to turn in my work, but it does say "if any" lol

OpenStudy (campbell_st):

well you could do a simple sketch... circle is centre (-2, 0) radius 2 and the line y = x - 4 has intercepts x = 4, y = -4  so all I would do is perhaps include a sketch simply what you have to \[x^2 - 2x + 8 = 0\] then say there are no points on intersection

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (campbell_st):

glad to help

OpenStudy (anonymous):

can you help with another? I have half of it I think.... @campbell_st

OpenStudy (campbell_st):

ok... just post it

OpenStudy (anonymous):

okay. \[2x+y=15\] \[(x-2)^{2}+(y-1)^{2}=25\] I know I have to get the y on alone, so i subtracted the 2x from each side and got \[y=-2x+15\] then i plugged that into the other equation and i guess that's where i screwed up. @campbell_st

OpenStudy (campbell_st):

I just graphed them again, and they don't intersect... circle has a centre at (2, 1) with radius 5 line has intercepts: x intercept 7.5 and y intercept 15 so no solutions... perhaps before you start use this site to graph them https://www.desmos.com/calculator

OpenStudy (anonymous):

geez... i wonder why they're on here... Thank you though.

OpenStudy (campbell_st):

lol... its a bit odd.... but I always sketch things 1st... it may help to make sense of the solutions...